OpenStudy (anonymous):

at 9:00 on a saturday morning, 2 bicyclists heading in opposite directions pass each other on a bicycle path. the bicyclist heading north is riding 5km / hour faster than the bicyclist heading south. at 10:15, they are 41.25 km apart. find the 2 bicyclists' rates. (please show work, i want to understand how you got the answer)

OpenStudy (anonymous):

r t = d Let r be the speed of the bicyclist heading South. Solve the following for r. (r + 5) 1.25 + r 1.25 = 41.25 r = 14 km/hr, r+5 = 19 km/hr

OpenStudy (anonymous):

its okay , at least you tried! better than what i came up with! :P it took me almost 30 minutes just to find the actual answer!

hero (hero):

Trying isn't good enough

hero (hero):

I realize what went wrong now. They had already given the total distance and total time

hero (hero):

I stupidly added them again

OpenStudy (anonymous):

relax, we're humans! we all make mistakes (:

hero (hero):

Weird thing is, I still keep getting the same thing

hero (hero):

Even after correcting the numbers

OpenStudy (anonymous):

all i know is that the answer is : the north bicyclist - 19 km & south bicyclist - 14 km. my teacher said the answers. all we have to do is figure out how thats the answer and show the work.

hero (hero):

will someone please explain to me my mistake?

hero (hero):

There's just no consistency with these Apparently: \[\frac{41.25}{1.25} = 2x + 5\] is the right setup to get those answers. I just don't get why you're not supposed to divide the rate by 2 to get the average rate.

hero (hero):

Lana, any sweet words of advice?

hero (hero):

I'm going to find out what the right way to setup the chart is. I don't give up on things until I figure out what the problem is

hero (hero):

The only thing I can't figure out is why I'm not supposed to find the average rate by dividing 2x+5 by 2

OpenStudy (anonymous):

divide by two, because of how many bicyclist are in the problem.... i think ?

hero (hero):

No, division by two is related to finding the average value. But it's not required to use in this case.